Assume that you have collected 35 ml egg white. Answer the following question: What will be the pH of the egg white solution after adding 1 N acetic acid, assuming egg white is neutral?
Answer #1

Let the volume of 1 N acetic acid, be V.

acetic acid Ka = 1.8 x 10^-5

[ H+ ] = sqrt(Ka x c)

We need concentration c .

3.5mL of 1N CH3COOH is added.

So Number of normals = 3.5 x 1 milli

Net vol = 38.5 ml

Normality = 3.5 * 1 /38.5 = 0.091 N

[ H+ ] = sqrt(Ka x c)

pH = -log( sqrt(Ka *c) ) = 2.89

pH = 2.89